MOA-II Obsevation Log
Return to index
Previous day
Next day
DATE: 2007/03/22
OBSERVER: Paul
CONDITION: Mostly clear
COMMENT: Warm camera, Camera Error at start of night. Tel driver unit error at start(Zero search). D2169(60), D2170(300)
Compressor gas pressure: 1.9 MPa
Expected Observation Time: START 19:55 END 05:56
TARGET | RUN | TOTAL RUN | TOTAL FIELD | TOTAL EXPOSURE (sec) |
GB | 5290-5323 | 34 | 192 | 11520 |
LMC | 1378-1381 | 4 | 19 | 5700 |
SMC | - | 0 | 0 | 0 |
ALERT | - | 0 | 0 | 0 |
GRB | - | 0 | 0 | 0 |
TOTAL | --- | 38 | 211 | 17220 (04h47m00s) |
GB FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
8 | 8 | 8 | 7 | 33 | 6 | 7 | 7 | 32 | 6 | 7 | 7 |
Field13 | Field14 | Field15 | Field16 | Field17 | Field18 | Field19 | Field20 | Field21 | Field22 | Field23 |
7 | 6 | 6 | 7 | 5 | 5 | 5 | 5 | 5 | 5 | 0 |
LMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
1 | 1 | 1 | 0 | 0 | 1 | 4 | 3 | 3 | 1 | 1 | 1 |
Field13 | Field14 | Field15 | Field16 |
1 | 1 | 0 | 0 |
SMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
SPECIAL TARGET NAME
CHECK SHEET
Time | Status | Weather |
22:20 | observing | clear |
23:00 | observing | clear |
00:00 | observing | clear |
01:00 | observing | clear |
02:00 | observing | clear |
03:00 | observing | clear |
04:00 | observing | clear |
05:00 | observing | maybe a few clouds |
CCD TEMPERATURE
Previous day
Next day

VACUUM
Previous day
Next day

TELESCOPE TEMPERATURE
Previous day
Next day

Return to index
Previous day
Next day